Cutter Enterprises purchased equipment for $72,000 on January 1, 2021. The equipment is expected to have a five-year life and a residual value of $6,000.

Using the sum-of-the-years'-digits method, depreciation for 2022 and book value at December 31, 2022, would be: (Do not round depreciation rate per year)

Multiple Choice

  • $19,200 and $30,800 respectively.

  • $17,600 and $26,400 respectively.

  • $19,200 and $28,800 respectively.

  • $17,600 and $32,400 respectively.
Answer #1

Depreciation for 2021 = (72,000-6,000)*5/15 = 22,000

Depreciation for 2022 = (72,000-6,000)*4/15 = 17,600

Book value = 72,000-22,000-17,600 = 32,400

Option D is the answer